OpenStudy (anonymous):

4(x+4)=6(x-2) solution set

OpenStudy (texaschic101):

4(x + 4) = 6(x - 2) -- distribute through the parenthesis 4x + 16 = 6x - 12 -- subtract 6x from both sides 4x - 6x + 16 = -12 -- subtract 16 from both sides 4x - 6x = -12 - 16 -- combine like terms -2x = -28 -- divide both sides by -2 x = -28/-2 x = 14 check... 4(14 + 4) = 6(14 - 2) 4(18) = 6(12) 72 = 72 (correct) x = 14 any questions ?